Travelling downstream, a boat can go 54 km in 3 hours. Going upstream, it takes twice as long to travel only 60 km. What is the speed of the boat in still water and what is the speed of the current?
Answer by TimothyLamb(4379)

s = d / t
s + c = 54/3
s - c = 60/6 = 10
s = 10 + c
---
s + c = 54/3
10 + c + c = 54/3
2c = 54/3 - 30/3
2c = 24/3 = 8
Answer:
c = 4 kph
s = 14 kph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php